1-1 12) Solve the given initial-value problem. 6 9 X = X, X(0) = X(t) =
Answer #1

A- 6 A-d I z 16-17 Lo -1 el -1 12-1 LA-AI1 (6-X) (12-1) + 72-122-61 + x² + 9 I X²-187 +81=0 (1-92=0 -3 Lo =) 1 =9, 6-9 9 -1 1

1) -1 8 [347 36-62 CI + (=8 3G-G2 = -1 3(82-C2 = -1 24-C2 = -1 C 24+1 = 25 gt X(t) = 8 e 3 + 25 te** [?]:est gt X(t): 75te gt